# TabImpute: Accurate and Fast Zero-Shot Missing-Data Imputation with a Pre-Trained Transformer

### v3

@@ 5 Conclusion & Future Work
In this paper, we present a state-of-the-art pre-trained transformer for the tabular missing data problem. We build on recent work in tabular representation learning by adapting TabPFN’s architecture and training pipeline for the missing data setting. Even though we train purely on synthetic data, we are able to impute entries accurately on real-world OpenML data for a comprehensive set of missingness patterns, thus showcasing our model’s ability to generalize to unseen domains. We open-source not only our model architecture and weights, but also our training and evaluation code (available at https://anonymous.4open.science/r/tabular-6F65/README.md). We hope this will facilitate others to validate, utilize, and build upon our work.
Since we use the same architecture as TabPFN, we also suffer from the quadratic time complexity of attention along both rows and columns. Due to entry-wise featurization, this complexity is squared again along the row axis. While TabImpute proved to be fast on relatively small tables in MissBench, we expect scalability to be a concern on larger table sizes. Additionally, TabImpute is very fast when using a GPU, but is slower on a CPU similar to TabPFN. This can be partly alleviated by imputing tables in chunks, but we leave this for future work.
Note that since we utilize a base architecture similar to TabPFN, any further improvements to TabPFN’s architecture can be immediately ported to TabImpute. For example, zeng2025tabflex and qu2025tabicl propose different attention mechanisms to speed up TabPFN-like architectures for tabular data. Finally, since we use a PFN architecture, we output a distribution for each missing entry and can sample from this distribution for multiple imputation (rubin2018multiple). In future work, we plan on (i) exploring training further on more complex missingness patterns and data-generating processes, (ii) enhancing our method to impute categorical data better, (iii) extending our evaluation set to causal inference settings, which can be modeled as missing-data problems (agarwal2023causal), (iv) improving the architecture to scale to larger datasets, and (v) utilizing TabImpute further for multiple imputation datasets.
Reproducibility Statement.
We open-source our training code, evaluation code, model architecture and weights, and synthetic data-generation pipeline (available at https://anonymous.4open.science/r/tabular-6F65/). While we utilize 8 H200 GPUs to train our model, our model should be able to fit easily on GPUs with far less memory because our model only has around 22 million parameters. All of the datasets we use for evaluation are public and available in OpenML, making it straightforward to reproduce our results from our pre-trained model.
@@ A.1 Additional tests
We support imputing categorical variables via one-hot encodings: First, we convert each categorical column into several one-hot encoding columns. Then, we impute missing entries within this now purely numerical matrix. We then choose the class with the highest score as the categorical imputation for that missing entry. Note that we could also perform a softmax operation over the imputed scores within the one-hot encoded columns to output probabilities over classes. With this method, TabImpute achieves AUC close to that of MissForest and HyperImpute when tested on MCAR missingness with $p=0.4$, as shown in Tab. 4. We leave it for future work to incorporate categorical features into the training procedure, which would likely improve TabImpute’s performance.
Table 4: Area Under the Curve (AUC) Performance on Categorical Columns with MCAR ($p=0.4$) missingness.
@@ A.2.2 Details on Seq-MNAR
This pattern simulates a scenario where masking matrix values $M_{ij}\in\{0,1\}$ for each column $j$ are adaptively chosen depending on the information up to column j1j-1 (i.e., regard columns as time). Specifically, we employ variants of bandit algorithms lattimore2020bandit while regarding the binary masking matrix values as the two arms. Such patterns commonly arise in sequential experiments ghosh2024did.
@@
We generate exogenous Gaussian noise and add it to the true matrix $X^{*}$ and regard $X^{*}$ as the reward for arm 0 and its noisy version as the reward for arm $1$. Then, starting from the first column with multiple rows as multiple agents, we randomly initiate (with random configurations) one of the four algorithms lattimore2020bandit: $\varepsilon$-greedy, Upper Confidence Bound (UCB) (auer2002finite), Thompson sampling with random configurations (thompson1933likelihood). Further, we have the option to randomly mix pooling techniques ghosh2024did on top of any of the four algorithms.
A.2.3 Self-Masking-MNAR
@@ A.2.6 Polarization-MNAR
This pattern simulates data where values falling in the middle of a feature’s distribution are preferentially removed, simulating survey non-response from individuals with moderate opinions. This is implemented by setting values between the $q$-th and (1q)(1-q)-th quantiles to missing. A ”soft” version makes the observation probability proportional to the value’s distance from the median. Polarization-MNAR is recognized in the survey methodology literature as a form of voluntary response bias or self-selection bias (bethlehem2010selection). One obvious example can be found in online product reviews, which frequently exhibit a U-shaped or J-shaped distribution, heavily skewed towards the highest and lowest ratings while the middle ground remains sparse—sometimes referred to as the ”brag-and-moan” bias (hu2009overcoming). Similarly, in political polling, nonresponse bias can be exacerbated if highly polarized individuals are more motivated to participate than moderates, potentially exaggerating measures of mass polarization cavari2023survey.
@@ A.2.9 Two-Phase-MNAR
A.3 Additional tables
@@
Column-wise mean (hawthorne2005imputing) Mean of columns
SoftImpute (hastie2015matrix) Iterative soft thresholding singular value decomposition based on a low-rank assumption on the data
$k$-Nearest Neighbors (60c19788-1128-3b5f-9275-2d63cc155832) Row-wise nearest neighbors mean
HyperImpute (jarrett2022hyperimpute) Iterative imputation method optimizing over a suite of imputation methods
Optimal transport method (muzellec2020missing) Uses optimal transport distances as a loss to impute missing values based on the principle that two randomly drawn batches from the same dataset should share similar data distributions
MissForest (missforest2011) Repeatedly trains a random forest model for each variable on the observed values to predict and fill in missing entries until convergence
ICE (ice2011) Imputation with iterative and chained equations of linear/logistic models for conditional expectations
MICE (royston2011multiple) Handles missing data by iteratively imputing each incomplete variable using regression models that condition on all other variables
GAIN (gain2018) Adapts generative adversarial networks (goodfellow2020generative) where the generator imputes missing values and the discriminator identifies which components are observed versus imputed
MIWAE (mattei2019miwae) Learns a deep latent variable model and then performs importance sampling for imputation
ForestDiffusion (jolicoeur2024generating) Trains a diffusion model using XGBoost directly on incomplete tabular data and then fills in missing values with an adapted inpainting algorithm
TabPFN The tabpfn-extensions package includes a part to impute missing entries in a table column-by-column using TabPFN.
@@
Pattern TabImpute EWF-TabPFN HyperImpute OT MissForest
MCAR 0.882 ± 0.021 0.817 ± 0.024 0.804 ± 0.031 0.855 ± 0.019 0.867 ± 0.022
NN-MNAR 0.900 ± 0.019 0.876 ± 0.020 0.760 ± 0.036 0.814 ± 0.021 0.818 ± 0.029
Self-Masking-MNAR 0.703 ± 0.038 0.682 ± 0.042 0.689 ± 0.039 0.568 ± 0.037 0.638 ± 0.040
Col-MAR 0.877 ± 0.030 0.855 ± 0.027 0.814 ± 0.040 0.750 ± 0.040 0.773 ± 0.039
Block-MNAR 0.890 ± 0.026 0.905 ± 0.026 0.873 ± 0.027 0.857 ± 0.024 0.860 ± 0.023
Seq-MNAR 0.920 ± 0.014 0.905 ± 0.015 0.864 ± 0.030 0.894 ± 0.014 0.829 ± 0.031
Panel-MNAR 0.915 ± 0.025 0.711 ± 0.048 0.865 ± 0.033 0.887 ± 0.025 0.912 ± 0.020
Polarization-MNAR 0.804 ± 0.024 0.879 ± 0.021 0.622 ± 0.039 0.802 ± 0.020 0.567 ± 0.033
Soft-Polarization-MNAR 0.759 ± 0.045 0.864 ± 0.024 0.711 ± 0.034 0.750 ± 0.029 0.667 ± 0.041
Latent-Factor-MNAR 0.891 ± 0.021 0.883 ± 0.018 0.775 ± 0.036 0.842 ± 0.026 0.834 ± 0.023
Cluster-MNAR 0.903 ± 0.014 0.871 ± 0.021 0.839 ± 0.028 0.828 ± 0.022 0.830 ± 0.021
Two-Phase-MNAR 0.886 ± 0.026 0.851 ± 0.029 0.865 ± 0.031 0.807 ± 0.028 0.873 ± 0.020
Censoring-MNAR 0.710 ± 0.032 0.593 ± 0.045 0.786 ± 0.041 0.589 ± 0.036 0.671 ± 0.038
Overall 0.849 ± 0.008 0.823 ± 0.009 0.790 ± 0.010 0.788 ± 0.009 0.780 ± 0.009
Pattern K-Nearest Neighbors ICE SoftImpute Col Mean DiffPuter
MCAR 0.750 ± 0.025 0.659 ± 0.041 0.617 ± 0.044 0.478 ± 0.041 0.478 ± 0.041
NN-MNAR 0.741 ± 0.026 0.627 ± 0.039 0.679 ± 0.035 0.497 ± 0.041 0.497 ± 0.041
Self-Masking-MNAR 0.614 ± 0.038 0.673 ± 0.049 0.396 ± 0.044 0.293 ± 0.038 0.293 ± 0.038
Col-MAR 0.839 ± 0.028 0.825 ± 0.038 0.703 ± 0.040 0.622 ± 0.046 0.598 ± 0.047
Block-MNAR 0.860 ± 0.028 0.812 ± 0.037 0.799 ± 0.042 0.782 ± 0.034 0.806 ± 0.029
Seq-MNAR 0.885 ± 0.018 0.815 ± 0.037 0.783 ± 0.036 0.775 ± 0.033 0.775 ± 0.033
Panel-MNAR 0.899 ± 0.028 0.797 ± 0.044 0.340 ± 0.053 0.787 ± 0.037 0.811 ± 0.032
Polarization-MNAR 0.641 ± 0.028 0.560 ± 0.041 0.665 ± 0.048 0.964 ± 0.012 0.964 ± 0.012
Soft-Polarization-MNAR 0.580 ± 0.042 0.729 ± 0.031 0.677 ± 0.033 0.747 ± 0.030 0.747 ± 0.030
Latent-Factor-MNAR 0.764 ± 0.029 0.703 ± 0.042 0.714 ± 0.039 0.590 ± 0.045 0.590 ± 0.045
Cluster-MNAR 0.774 ± 0.024 0.750 ± 0.042 0.627 ± 0.046 0.558 ± 0.044 0.558 ± 0.044
Two-Phase-MNAR 0.898 ± 0.020 0.876 ± 0.032 0.678 ± 0.047 0.671 ± 0.042 0.671 ± 0.042
Censoring-MNAR 0.515 ± 0.047 0.866 ± 0.035 0.455 ± 0.054 0.346 ± 0.040 0.307 ± 0.037
Overall 0.751 ± 0.010 0.746 ± 0.011 0.626 ± 0.013 0.624 ± 0.013 0.623 ± 0.013
Pattern TabPFN MICE MIWAE GAIN
MCAR 0.415 ± 0.041 0.329 ± 0.044 0.254 ± 0.043 0.545 ± 0.039
NN-MNAR 0.428 ± 0.041 0.300 ± 0.049 0.230 ± 0.043 0.392 ± 0.049
Self-Masking-MNAR 0.268 ± 0.037 0.611 ± 0.051 0.217 ± 0.041 0.452 ± 0.056
Col-MAR 0.549 ± 0.050 0.520 ± 0.054 0.488 ± 0.048 0.275 ± 0.054
Block-MNAR 0.582 ± 0.051 0.502 ± 0.048 0.606 ± 0.039 0.202 ± 0.051
Seq-MNAR 0.638 ± 0.045 0.475 ± 0.050 0.584 ± 0.045 0.204 ± 0.048
Panel-MNAR 0.234 ± 0.042 0.553 ± 0.048 0.269 ± 0.045 0.536 ± 0.059
Polarization-MNAR 0.907 ± 0.029 0.192 ± 0.037 0.562 ± 0.027 0.294 ± 0.045
Soft-Polarization-MNAR 0.600 ± 0.050 0.219 ± 0.038 0.548 ± 0.036 0.425 ± 0.062
Latent-Factor-MNAR 0.499 ± 0.048 0.338 ± 0.047 0.355 ± 0.045 0.296 ± 0.046
Cluster-MNAR 0.499 ± 0.047 0.327 ± 0.044 0.330 ± 0.044 0.357 ± 0.048
Two-Phase-MNAR 0.478 ± 0.049 0.564 ± 0.048 0.529 ± 0.049 0.228 ± 0.049
Censoring-MNAR 0.342 ± 0.041 0.648 ± 0.050 0.308 ± 0.045 0.320 ± 0.051
Overall 0.495 ± 0.014 0.429 ± 0.014 0.407 ± 0.013 0.348 ± 0.015
@@
Pattern ICE HyperImpute MissForest K-Nearest Neighbors MICE
MCAR 0.940 ± 0.013 0.919 ± 0.019 0.825 ± 0.019 0.742 ± 0.017 0.814 ± 0.033
NN-MNAR 0.918 ± 0.019 0.925 ± 0.020 0.852 ± 0.020 0.723 ± 0.019 0.819 ± 0.030
Self-Masking-MNAR 0.788 ± 0.035 0.719 ± 0.037 0.641 ± 0.036 0.606 ± 0.034 0.727 ± 0.054
Seq-MNAR 0.827 ± 0.025 0.842 ± 0.020 0.814 ± 0.019 0.785 ± 0.021 0.853 ± 0.030
Polarization-MNAR 0.616 ± 0.039 0.648 ± 0.042 0.540 ± 0.039 0.718 ± 0.032 0.300 ± 0.041
Soft-Polarization-MNAR 0.787 ± 0.038 0.757 ± 0.039 0.750 ± 0.045 0.529 ± 0.046 0.434 ± 0.043
Latent-Factor-MNAR 0.930 ± 0.016 0.912 ± 0.023 0.862 ± 0.019 0.735 ± 0.018 0.821 ± 0.032
Cluster-MNAR 0.919 ± 0.018 0.913 ± 0.021 0.830 ± 0.020 0.736 ± 0.017 0.811 ± 0.037
Two-Phase-MNAR 0.831 ± 0.030 0.795 ± 0.030 0.758 ± 0.023 0.835 ± 0.021 0.771 ± 0.041
Overall 0.840 ± 0.010 0.826 ± 0.011 0.764 ± 0.011 0.712 ± 0.010 0.706 ± 0.016
Pattern ForestDiffusion OT TabImpute EWF-TabPFN SoftImpute
MCAR 0.822 ± 0.021 0.721 ± 0.017 0.724 ± 0.029 0.520 ± 0.035 0.465 ± 0.060
NN-MNAR 0.756 ± 0.017 0.697 ± 0.018 0.695 ± 0.031 0.541 ± 0.034 0.400 ± 0.064
Self-Masking-MNAR 0.673 ± 0.035 0.528 ± 0.029 0.640 ± 0.035 0.601 ± 0.039 0.364 ± 0.048
Seq-MNAR 0.905 ± 0.018 0.739 ± 0.019 0.621 ± 0.029 0.576 ± 0.038 0.624 ± 0.048
Polarization-MNAR 0.267 ± 0.045 0.865 ± 0.024 0.866 ± 0.023 0.862 ± 0.027 0.640 ± 0.041
Soft-Polarization-MNAR 0.507 ± 0.036 0.662 ± 0.036 0.632 ± 0.040 0.747 ± 0.034 0.561 ± 0.042
Latent-Factor-MNAR 0.747 ± 0.022 0.707 ± 0.019 0.676 ± 0.032 0.529 ± 0.040 0.459 ± 0.059
Cluster-MNAR 0.757 ± 0.024 0.704 ± 0.018 0.663 ± 0.031 0.548 ± 0.040 0.397 ± 0.063
Two-Phase-MNAR 0.909 ± 0.025 0.678 ± 0.029 0.742 ± 0.029 0.643 ± 0.040 0.597 ± 0.046
Overall 0.705 ± 0.014 0.700 ± 0.009 0.695 ± 0.011 0.619 ± 0.013 0.501 ± 0.018
Pattern MIWAE GAIN Col Mean TabPFN
MCAR 0.420 ± 0.018 0.600 ± 0.034 0.024 ± 0.008 0.067 ± 0.016
NN-MNAR 0.414 ± 0.014 0.491 ± 0.036 0.031 ± 0.015 0.040 ± 0.013
Self-Masking-MNAR 0.225 ± 0.037 0.463 ± 0.057 0.143 ± 0.030 0.156 ± 0.031
Seq-MNAR 0.589 ± 0.026 0.218 ± 0.043 0.282 ± 0.041 0.285 ± 0.038
Polarization-MNAR 0.682 ± 0.033 0.318 ± 0.047 0.783 ± 0.027 0.741 ± 0.034
Soft-Polarization-MNAR 0.532 ± 0.040 0.391 ± 0.065 0.436 ± 0.033 0.331 ± 0.039
Latent-Factor-MNAR 0.468 ± 0.024 0.361 ± 0.042 0.106 ± 0.028 0.135 ± 0.027
Cluster-MNAR 0.452 ± 0.021 0.381 ± 0.042 0.059 ± 0.020 0.102 ± 0.024
Two-Phase-MNAR 0.555 ± 0.039 0.169 ± 0.043 0.409 ± 0.041 0.398 ± 0.039
Overall 0.482 ± 0.012 0.377 ± 0.017 0.252 ± 0.016 0.250 ± 0.015
Table 8: Mean Column-wise $R^{2}$ ± Standard Error by Missingness Pattern

### v4

@@ 5 Conclusion & Future Work
In this paper, we present TabImpute, a pre-trained transformer for imputing missing data in small tables. We build on recent work in tabular representation learning by adapting TabPFN’s architecture and training pipeline for the missing data setting. Though TabImpute is trained purely on synthetic data, it accurately imputes entries on real-world OpenML data across a comprehensive set of missingness patterns, demonstrating its ability to generalize to unseen domains. We open-source not only our model architecture and weights, but also our training and evaluation code. We hope this will facilitate others in validating, utilizing, and building upon our work.
While TabImpute performs well on small tables, it is not designed for very large tables because of the entry-wise featurization and quadratic attention cost. However, since we use the same network architecture as TabPFN, any further improvements to TabPFN’s architecture can be ported to TabImpute immediately. For example, (zeng2025tabflex) and (qu2025tabicl) propose different attention mechanisms to speed up TabPFN-like architectures for tabular data.
In future work, we plan on (i) exploring training further on more complex missingness patterns and data-generating processes, (ii) enhancing our method to impute categorical data better, (iii) extending our evaluation set to causal-inference settings, which can be modeled as missing-data problems (agarwal2023causal), and (iv) improving the architecture to scale to larger datasets.
Impact statement.
Our work seeks to advance the fields of machine learning and data science by improving missing-data imputation. While there are many possible impacts on society as a whole from our work, we do not feel that any direct impacts on society need to be highlighted here.
@@ A.1 Additional tests
We support imputing categorical variables via one-hot encodings: First, we convert each categorical column into several one-hot encoding columns. Then, we impute missing entries within this now purely numerical matrix. We then choose the class with the highest score as the categorical imputation for that missing entry. Note that we could also perform a softmax operation over the imputed scores within the one-hot encoded columns to output probabilities over classes. With this method, TabImpute achieves AUC close to that of MissForest and HyperImpute when tested on MCAR missingness with the probability of missing set to 0.4, as shown in Tab. 4. We leave incorporating categorical features into the pre-training procedure for future work.
Table 4: Area Under the Curve (AUC) Performance on Categorical Columns with MCAR (Probability of Missing = 0.4) missingness.
@@ A.2.2 Details on Seq-MNAR
This pattern simulates a scenario where masking matrix values $M_{ij}\in\{0,1\}$ for each column $j$ are adaptively chosen depending on the information up to column j1j-1 (i.e., regard columns as time). Specifically, we employ variants of bandit algorithms (lattimore2020bandit) while regarding the binary masking matrix values as the two arms. Such patterns commonly arise in sequential experiments (ghosh2024did).
@@
We generate exogenous Gaussian noise and add it to the true matrix $X^{*}$ and regard $X^{*}$ as the reward for arm 0 and its noisy version as the reward for arm $1$. Then, starting from the first column with multiple rows as multiple agents, we randomly initiate (with random configurations) one of the four algorithms (lattimore2020bandit): $\varepsilon$-greedy, Upper Confidence Bound (UCB) (auer2002finite), Thompson sampling with random configurations (thompson1933likelihood). Further, we have the option to randomly mix pooling techniques (ghosh2024did) on top of any of the four algorithms.
A.2.3 Self-MNAR
@@ A.2.6 Polarization-MNAR
This pattern simulates data where values falling in the middle of a feature’s distribution are preferentially removed, simulating survey non-response from individuals with moderate opinions. This is implemented by setting values between the $q$-th and (1q)(1-q)-th quantiles to missing. A ”soft” version makes the observation probability proportional to the value’s distance from the median. Polarization-MNAR is recognized in the survey methodology literature as a form of voluntary response bias or self-selection bias (bethlehem2010selection). One obvious example can be found in online product reviews, which frequently exhibit a U-shaped or J-shaped distribution, heavily skewed towards the highest and lowest ratings while the middle ground remains sparse—sometimes referred to as the ”brag-and-moan” bias (hu2009overcoming). Similarly, in political polling, nonresponse bias can be exacerbated if highly polarized individuals are more motivated to participate than moderates, potentially exaggerating measures of mass polarization (cavari2023survey).
@@ A.2.9 Two-Phase-MNAR
A.3 Ablation of Featurization and Pre-Training
[…]
A possible explanation for the effectiveness of entry-wise featurization is that it allows entry-specific information to route through the network without interference from other prediction targets, whereas Col-TabPFN entangles prediction routes. This role is analogous to how CLS tokens in BERT enable sentence-level information to be aggregated without interfering with token-level representations (devlin2019bert).
A.4 Additional tables
@@
SoftImpute (hastie2015matrix) Iterative soft thresholding singular value decomposition based on a low-rank assumption on the data.
$k$-Nearest Neighbors (60c19788-1128-3b5f-9275-2d63cc155832) Row-wise nearest neighbors mean. We set $k=5$ for our tests.
HyperImpute (jarrett2022hyperimpute) Iterative imputation method optimizing over a suite of imputation methods.
Optimal transport method (muzellec2020missing) Uses optimal transport distances as a loss to impute missing values based on the principle that two randomly drawn batches from the same dataset should share similar data distributions.
MissForest (missforest2011) Repeatedly trains a random forest model for each variable on the observed values to predict and fill in missing entries until convergence.
ICE (ice2011) Imputation with iterative and chained equations of linear/logistic models for conditional expectations.
MICE (royston2011multiple) Handles missing data by iteratively imputing each incomplete variable using regression models that condition on all other variables.
GAIN (gain2018) Adapts generative adversarial networks (goodfellow2020generative) where the generator imputes missing values and the discriminator identifies which components are observed versus imputed.
MIWAE (mattei2019miwae) Learns a deep latent variable model and then performs importance sampling for imputation.
ForestDiffusion (jolicoeur2024generating) Trains a diffusion model using XGBoost directly on incomplete tabular data and then fills in missing values with an adapted inpainting algorithm.
TabPFN (hollmann2023tabpfn) The tabpfn-extensions package includes a part to impute missing entries in a table column-by-column using TabPFN.
ReMasker (du2023remasker) A transformer-based method using masked autoencoding.
CACTI (gorla2025cacti) A transformer-based method that leverages copy-masking and leverages the text-based contextual information of column names. We use the version that does not use textual information (CMAE) to keep our tests fair.
@@
Pattern TabImpute EWF-TabPFN OT HyperImpute MissForest
MCAR 0.904 ± 0.018 0.826 ± 0.023 0.866 ± 0.019 0.814 ± 0.027 0.854 ± 0.025
NN-MNAR 0.943 ± 0.014 0.916 ± 0.016 0.873 ± 0.014 0.785 ± 0.035 0.838 ± 0.026
Self-MNAR 0.775 ± 0.030 0.765 ± 0.033 0.693 ± 0.033 0.751 ± 0.034 0.700 ± 0.034
Col-MAR 0.921 ± 0.018 0.888 ± 0.014 0.788 ± 0.032 0.868 ± 0.029 0.830 ± 0.025
Block-MNAR 0.937 ± 0.013 0.923 ± 0.014 0.857 ± 0.027 0.854 ± 0.027 0.850 ± 0.024
Seq-MNAR 0.930 ± 0.015 0.915 ± 0.013 0.853 ± 0.029 0.837 ± 0.031 0.802 ± 0.033
Panel-MNAR 0.955 ± 0.008 0.889 ± 0.019 0.918 ± 0.011 0.883 ± 0.031 0.897 ± 0.019
Polarization-MNAR 0.825 ± 0.025 0.896 ± 0.020 0.787 ± 0.021 0.631 ± 0.039 0.580 ± 0.033
Soft-Polarization-MNAR 0.805 ± 0.047 0.867 ± 0.022 0.773 ± 0.029 0.669 ± 0.036 0.625 ± 0.039
Latent-Factor-MNAR 0.923 ± 0.016 0.874 ± 0.021 0.870 ± 0.019 0.783 ± 0.038 0.835 ± 0.026
Cluster-MNAR 0.933 ± 0.011 0.893 ± 0.020 0.868 ± 0.016 0.844 ± 0.029 0.848 ± 0.021
Two-Phase-MNAR 0.914 ± 0.012 0.872 ± 0.023 0.822 ± 0.025 0.880 ± 0.029 0.881 ± 0.016
Censoring-MNAR 0.766 ± 0.030 0.660 ± 0.041 0.739 ± 0.021 0.810 ± 0.028 0.749 ± 0.023
Overall 0.887 ± 0.007 0.860 ± 0.007 0.824 ± 0.007 0.801 ± 0.009 0.791 ± 0.008
Pattern K-NN ICE CACTI SoftImpute ForestDiffusion
MCAR 0.772 ± 0.022 0.660 ± 0.042 0.714 ± 0.035 0.632 ± 0.044 0.450 ± 0.033
NN-MNAR 0.800 ± 0.020 0.665 ± 0.038 0.727 ± 0.035 0.734 ± 0.034 0.496 ± 0.036
Self-MNAR 0.680 ± 0.037 0.644 ± 0.052 0.652 ± 0.041 0.511 ± 0.044 0.662 ± 0.036
Col-MAR 0.872 ± 0.017 0.845 ± 0.034 0.806 ± 0.024 0.720 ± 0.030 0.511 ± 0.038
Block-MNAR 0.893 ± 0.014 0.806 ± 0.032 0.752 ± 0.045 0.850 ± 0.027 0.740 ± 0.030
Seq-MNAR 0.851 ± 0.027 0.784 ± 0.037 0.780 ± 0.033 0.758 ± 0.042 0.603 ± 0.030
Panel-MNAR 0.940 ± 0.011 0.819 ± 0.039 0.875 ± 0.025 0.671 ± 0.043 0.661 ± 0.040
Polarization-MNAR 0.658 ± 0.028 0.566 ± 0.041 0.511 ± 0.031 0.651 ± 0.048 0.291 ± 0.042
Soft-Polarization-MNAR 0.635 ± 0.035 0.681 ± 0.037 0.561 ± 0.040 0.719 ± 0.034 0.749 ± 0.033
Latent-Factor-MNAR 0.810 ± 0.022 0.719 ± 0.042 0.772 ± 0.027 0.724 ± 0.038 0.584 ± 0.030
Cluster-MNAR 0.807 ± 0.018 0.744 ± 0.044 0.742 ± 0.034 0.672 ± 0.045 0.579 ± 0.035
Two-Phase-MNAR 0.905 ± 0.015 0.875 ± 0.028 0.858 ± 0.031 0.691 ± 0.044 0.575 ± 0.040
Censoring-MNAR 0.561 ± 0.045 0.795 ± 0.036 0.721 ± 0.036 0.448 ± 0.049 0.738 ± 0.039
Overall 0.783 ± 0.009 0.739 ± 0.011 0.729 ± 0.010 0.675 ± 0.012 0.588 ± 0.011
Pattern ReMasker Col-TabPFN MIWAE MICE GAIN
MCAR 0.647 ± 0.032 0.448 ± 0.043 0.276 ± 0.046 0.350 ± 0.046 0.556 ± 0.040
NN-MNAR 0.397 ± 0.046 0.546 ± 0.043 0.366 ± 0.051 0.360 ± 0.050 0.507 ± 0.040
Self-MNAR 0.570 ± 0.055 0.396 ± 0.046 0.391 ± 0.045 0.534 ± 0.057 0.557 ± 0.054
Col-MAR 0.662 ± 0.046 0.575 ± 0.050 0.513 ± 0.047 0.477 ± 0.053 0.355 ± 0.057
Block-MNAR 0.685 ± 0.045 0.651 ± 0.046 0.661 ± 0.028 0.537 ± 0.038 0.217 ± 0.051
Seq-MNAR 0.540 ± 0.045 0.646 ± 0.046 0.552 ± 0.045 0.423 ± 0.048 0.258 ± 0.044
Panel-MNAR 0.681 ± 0.051 0.340 ± 0.049 0.498 ± 0.051 0.471 ± 0.048 0.493 ± 0.053
Polarization-MNAR 0.378 ± 0.043 0.908 ± 0.030 0.625 ± 0.027 0.224 ± 0.038 0.329 ± 0.042
Soft-Polarization-MNAR 0.417 ± 0.056 0.682 ± 0.043 0.605 ± 0.033 0.262 ± 0.039 0.502 ± 0.060
Latent-Factor-MNAR 0.462 ± 0.047 0.581 ± 0.049 0.428 ± 0.047 0.368 ± 0.051 0.386 ± 0.048
Cluster-MNAR 0.516 ± 0.053 0.596 ± 0.047 0.433 ± 0.049 0.356 ± 0.046 0.427 ± 0.046
Two-Phase-MNAR 0.736 ± 0.049 0.529 ± 0.047 0.538 ± 0.043 0.530 ± 0.049 0.249 ± 0.049
Censoring-MNAR 0.673 ± 0.050 0.397 ± 0.044 0.368 ± 0.047 0.554 ± 0.049 0.445 ± 0.054
Overall 0.567 ± 0.014 0.561 ± 0.014 0.481 ± 0.013 0.419 ± 0.014 0.406 ± 0.014
@@
Pattern ICE HyperImpute MissForest K-NN MICE
MCAR 0.941 ± 0.012 0.919 ± 0.019 0.823 ± 0.019 0.735 ± 0.018 0.812 ± 0.033
NN-MNAR 0.916 ± 0.020 0.922 ± 0.020 0.862 ± 0.020 0.740 ± 0.023 0.830 ± 0.030
Self-MNAR 0.790 ± 0.033 0.727 ± 0.036 0.655 ± 0.035 0.620 ± 0.033 0.741 ± 0.052
Seq-MNAR 0.833 ± 0.025 0.848 ± 0.020 0.818 ± 0.019 0.794 ± 0.021 0.854 ± 0.028
Polarization-MNAR 0.662 ± 0.037 0.689 ± 0.040 0.579 ± 0.041 0.744 ± 0.032 0.371 ± 0.047
Soft-Polarization-MNAR 0.809 ± 0.037 0.786 ± 0.039 0.783 ± 0.041 0.628 ± 0.042 0.551 ± 0.043
Latent-Factor-MNAR 0.934 ± 0.016 0.917 ± 0.022 0.875 ± 0.020 0.748 ± 0.021 0.833 ± 0.031
Cluster-MNAR 0.918 ± 0.019 0.913 ± 0.021 0.836 ± 0.021 0.756 ± 0.020 0.828 ± 0.035
Two-Phase-MNAR 0.832 ± 0.030 0.797 ± 0.030 0.760 ± 0.023 0.838 ± 0.019 0.769 ± 0.041
Overall 0.848 ± 0.010 0.835 ± 0.010 0.777 ± 0.010 0.734 ± 0.009 0.732 ± 0.015
Pattern ForestDiffusion OT TabImpute CACTI EWF-TabPFN
MCAR 0.819 ± 0.021 0.715 ± 0.017 0.720 ± 0.029 0.659 ± 0.037 0.508 ± 0.036
NN-MNAR 0.777 ± 0.018 0.724 ± 0.021 0.712 ± 0.033 0.707 ± 0.041 0.578 ± 0.038
Self-MNAR 0.679 ± 0.034 0.551 ± 0.030 0.656 ± 0.035 0.611 ± 0.040 0.625 ± 0.038
Seq-MNAR 0.909 ± 0.017 0.745 ± 0.020 0.635 ± 0.029 0.681 ± 0.035 0.588 ± 0.039
Polarization-MNAR 0.350 ± 0.048 0.876 ± 0.023 0.885 ± 0.019 0.592 ± 0.037 0.878 ± 0.025
Soft-Polarization-MNAR 0.597 ± 0.037 0.720 ± 0.036 0.675 ± 0.043 0.675 ± 0.040 0.781 ± 0.032
Latent-Factor-MNAR 0.765 ± 0.025 0.727 ± 0.022 0.695 ± 0.034 0.726 ± 0.031 0.552 ± 0.044
Cluster-MNAR 0.786 ± 0.024 0.728 ± 0.021 0.687 ± 0.034 0.732 ± 0.032 0.572 ± 0.045
Two-Phase-MNAR 0.907 ± 0.025 0.683 ± 0.028 0.746 ± 0.028 0.821 ± 0.035 0.643 ± 0.041
Overall 0.732 ± 0.013 0.719 ± 0.009 0.712 ± 0.011 0.689 ± 0.013 0.636 ± 0.014
Pattern ReMasker SoftImpute MIWAE GAIN Col-TabPFN
MCAR 0.805 ± 0.026 0.464 ± 0.058 0.407 ± 0.021 0.595 ± 0.033 0.049 ± 0.017
NN-MNAR 0.659 ± 0.055 0.479 ± 0.063 0.466 ± 0.029 0.554 ± 0.038 0.140 ± 0.041
Self-MNAR 0.501 ± 0.054 0.406 ± 0.047 0.283 ± 0.038 0.502 ± 0.054 0.214 ± 0.036
Seq-MNAR 0.515 ± 0.048 0.636 ± 0.048 0.596 ± 0.029 0.259 ± 0.041 0.302 ± 0.043
Polarization-MNAR 0.398 ± 0.044 0.675 ± 0.042 0.709 ± 0.034 0.392 ± 0.050 0.757 ± 0.034
Soft-Polarization-MNAR 0.297 ± 0.054 0.645 ± 0.040 0.613 ± 0.041 0.494 ± 0.061 0.453 ± 0.046
Latent-Factor-MNAR 0.634 ± 0.048 0.507 ± 0.060 0.502 ± 0.033 0.430 ± 0.044 0.198 ± 0.043
Cluster-MNAR 0.639 ± 0.054 0.436 ± 0.062 0.496 ± 0.031 0.457 ± 0.041 0.185 ± 0.043
Two-Phase-MNAR 0.656 ± 0.050 0.596 ± 0.046 0.563 ± 0.038 0.171 ± 0.044 0.408 ± 0.039
Overall 0.567 ± 0.018 0.538 ± 0.018 0.515 ± 0.013 0.428 ± 0.017 0.301 ± 0.016
Table 8: Mean Column-wise $R^{2}$ ± Standard Deviation by Missingness Pattern